Azimuth Fiducial Alignment
Fiducial blocks 1-7 were aligned then all 24 fiducials were tested by checking to see if the MCP received the 200ns interrupt pulse. We found that block number 15 and 16 did not interrupt the MCP so we realigned 16 but could not get 15 to work.
Rotator Read Heads
Read Heads #1 , 2 & 3 were aligned and tested by connecting each one to the MEI and running the rotator around in closed loop. Each one was then connected to Claudio’s fiducial board and checked to see if the interrupt pulse to the MEI board was there when a reference mark was passed. We did not check to see if every reference mark around the rotator produced an interrupt.
Rotator Limit Switches
The rotator 280 and 290 degree CW and CCW limit switches were checked to make sure they were operating properly. We found that the direction switch was backwards which made it trip on the wrong accuator. We also noticed that the bidirectional hard limit would trip the amplifier a couple of times and did not find the reason.
Rotator J.B. Cables
The serial cable and the power cable for the rotator J.B. were shortened to connect to the header plate, and 8-pin Burndy’s were installed because the hole in the header plate was drilled for 8-pin instead of 4-pin Burndy’s. The wrap cables also had the 8-pin Burndy’s installed.
The interlock display was checked and we found that Charlie needed to broadcast a new interlock packet because some of the bits had changed and new ones were added. The display was changed so that the whole display could be seen on the monitor in the lower level of the telescope (it runs at 1024 x 768). Craig thought that it should be optimized for that resolution so it could also be seen on some of the laptops. More work has to be done with the logic ladder to get it working properly.
Junction Box for Primary Mirror Temp Sensors
A location for the J.B. was discussed with Dan and it looks like the best spot would be the back of the PSS by the dog house. Talked to Steve about mounting it there so the next time out we could start installing the new Flex I/O modules.
Slip Detection Encoder
The slip detection encoder for the rotator motor was wired up to the Burndy on the encoder housing so Steve could mount the encoder.
Azimuth Stow Switch
The Front(East) azimuth stow switch was stuck in the compressed position which caused the stow light to come on twice 6 degrees apart. The switch was removed and a burr was removed and a lubricant was added. It was then remounted but not aligned. Correct alignment of the stow switch must still be checked.
Removed the 4 Interpolators for therotator and checked the switch settings which were S1-S3, S5, S8 off, S4, S6, S7 on. Reinstalled them back into the PSS.
Last modified 03/08/99